OpenStudy (anonymous):

What is the limit of s as it approaches 7? lim [(1/s)-(1/7)]/s-7

terenzreignz (terenzreignz):

\[\Large \lim_{s\rightarrow7}\left[\frac1s - \frac17\right](s-7)\]

terenzreignz (terenzreignz):

Did you try simplifying?

terenzreignz (terenzreignz):

Or actually, never mind... why don't you just downright replace s with 7? It doesn't result in anything ugly or undefined or anything.

OpenStudy (anonymous):

its 1/s - 1/7 OVER s-7

terenzreignz (terenzreignz):

Oh... sorry :) I just had that feeling that it was too easy.. but anyway \[\huge\lim_{s\rightarrow 7}\frac{\frac1s - \frac17}{s-7}\]

OpenStudy (anonymous):

yup!

terenzreignz (terenzreignz):

Well, I suggest you simplify \[\Large \frac1s - \frac17\] into just one fraction.

OpenStudy (anonymous):

but wouldn't equal to 0?

terenzreignz (terenzreignz):

It won't. Anyway, simplify?

OpenStudy (anonymous):

okey how do you do that?

terenzreignz (terenzreignz):

Seriously? -_- That's elementary stuff.... and you're in calculus XD But... not here to judge (not so harshly anyway :P ) \[\Large \frac{a}b-\frac{c}d = \frac{ad - bc}{bd}\]

terenzreignz (terenzreignz):

Stay with me here @jbeloy17

OpenStudy (anonymous):

sorry about that.

OpenStudy (anonymous):

i think i got it.. is the answer 51/2500?